Mechanics

A particle of mass 2kg located at the position (i+j) m has a velocity 2(+i-j+k) m/s . Its angular momentum about z-axis in kg-m2/s isa) 0b) +8c) 12d) -8

Aakash Gupta
8 Years agoGrade

1 Answer

Arun
8 Years ago
Dear Aakash
 
OP = sqrt(2)
 
Now
 
Angular momentum L = - m* Vxy * OP
L = - 2 * 2 sqrt(2) * sqrt(2)
 
L = - 8 kg m²/sec
 
 
Regards
Arun (askIITians forum expert)